OpenStudy (abbles):

What is the simplified form of: (x/3 - y/4)/(x/4 + y/3)

OpenStudy (anonymous):

\[=\frac{ x }{3 }*\frac{ x }{4 }+\frac{ x }{3 }*\frac{ y }{3 }-\frac{ x }{ 4 }*\frac{ y }{ 4 }-\frac{ y }{4 }*\frac{ y }{ 3 }\] \[=\frac{ x ^{2} }{12 }+\frac{ xy }{ 9 }-\frac{ xy }{ 16 }-\frac{ y ^{2} }{12 }\] simplify it .